OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I'll fan and give a medal!!

OpenStudy (abbycross167):

A line with slope 1/3 passes through the point (3, 6). Which point also lies on this line? a. (6, 5) b. (7, 6) c. (-3, -3) d. (-6, 3)

OpenStudy (abbycross167):

@Abbles could you please try to help me?

OpenStudy (robtobey2):

\[y=\frac{x}{3}+5 \]d.

OpenStudy (asiannerd):

Try to graph it. *the answer is on the negative side*

OpenStudy (pythagoras123):

General equation of a line: \[y-y _{1}=m(x-x _{1})\] where m is the gradient. Your equation is thus y-6=1/3(x-3) \[y=\frac{ 1 }{ 3 } x +5. \] Sub each given point in the equation and equate both sides of the equation to confirm if the given point satisfies the equation. If it does, then the point lies on the line.